OpenStudy (coconut):

What is the sum of the geometric sequence 4, 12, 36 … if there are 9 terms?

OpenStudy (anonymous):

\[\LARGE S_9=a_1\cdot \frac{1-q^9}{1-q}\] and... if a_1=4 ,a_2=12 \[\LARGE q={a_2 \over a_1}={12 \over 4}=3\] so q=3 thus... \[\LARGE S_9=4\cdot \frac{1-3^9}{1-3}\] I think this should be :)

OpenStudy (coconut):

is it in the formula a1(1-r^n) / 1-r

OpenStudy (anonymous):

this: \[\LARGE Sn=a_1\cdot {1-r^n \over 1-r}\] it's the same as: \[\LARGE Sn=a_1\cdot {1-q^n \over 1-q}\] just different letters :) ...

OpenStudy (coconut):

oh okay :) thank you! so would it be 8(1-(19683) / -2

OpenStudy (anonymous):

... quite close :) it's: \[\LARGE S_n=a_1\cdot {1-r^n \over 1-r}\] \[\LARGE S_9=4\cdot {1-3^9 \over 1-3}\] \[\LARGE S_9={4(1-19683) \over -2 }\] Rembember \[\LARGE a_1=4 \] :)

OpenStudy (coconut):

would it be 39364? thats what i got? :)